A father is four times as old as his daughter. In 6 years, he will be three times as old as she is. How old is the daughter now?
Answer by Maths68(1474)


Let
Present age of father = f
Present age of daughter = d

Given
=====
father is four times as old as his daughter
f=4d.............(1)
In 6 years
father will be = f+6
daughter will be = d+6
Then
father will be three times as old as daughter
f+6=3(d+6)
f+6=3d+18
f=3d+18-6
f=3d+12...............(2)
Put the value of f from (1) to (2)
f=3d+12...............(2)
4d=3d+12
4d-3d=12
d=12
Put the value of d in (1)
f=4d
f=4(12)
f=48



Present age of father = f = 48 years old
Present age of daughter = d = 12 years old